# Resolving phonon-mediated superconducting pairing symmetries from first-principles calculation

### v2

@@ II.3 DDM solution
$|\sigma_{i}\rangle$ $=$ $[\xi_{i}(n_{1}\mathbf{k}_{1}),\xi_{i}(n_{1}\mathbf{k}_{2}),...]^{T},$ (16)
$|\tau_{i}\rangle$ $=$ $[\xi_{i}(T_{0},n_{1}\mathbf{k}_{1}),\xi_{i}(T_{+},n_{1}\mathbf{k}_{1}),\xi_{i}(T_{-},n_{1}\mathbf{k}_{1}),...]^{T}$
reflect the $\mathbf{k}$-space symmetry of the corresponding pairing channel, which can be classified by the irreducible representations (irreps.) of the symmetry group.
There is still one remaining issue. So far, $\hat{\Delta}(n\mathbf{k}s_{1}s_{2})=\tilde{c}_{n,\mathbf{-k},s_{2}}c_{n\mathbf{k}s_{1}}$ and $\hat{\Delta}(n,-\mathbf{k}s_{2}s_{1})=\tilde{c}_{n\mathbf{k}s_{1}}c_{n,\mathbf{-k},s_{2}}$ are treated as independent operators, which are actually related by an undecided gauge transformation and the fermion anticommutation relation. (Mathematically, the basis of $\mathbb{V}$ is overcomplete.) Therefore, additional constraints should be applied to the eigenvectors.
Specifically, the symmetrized even-parity operator $\hat{\Delta}^{S}(n\mathbf{k})$ is U(1) and SU(2) invariant. The constraint to $|\sigma_{i}\rangle$ is simply $\xi_{i}(n_{m}\mathbf{k}_{j})=\xi_{i}(n_{m},-\mathbf{k}_{j})$. In practice, it means that eigenvectors inconsistent with it should be discarded.
For the odd-parity sector, the eigenvector coefficients $\xi_{i}(T_{0},n_{1}\mathbf{k}_{1})$, $\xi_{i}(T_{+},n_{1}\mathbf{k}_{1})$ and $\xi_{i}(T_{-},n_{1}\mathbf{k}_{1})$ mutually transform under the SU(2) rotation like a spin-1 object. Our recipe is to find a rotation such that $\xi_{i}(T_{+},n_{1}\mathbf{k}_{1})$ and $\xi_{i}(T_{-},n_{1}\mathbf{k}_{1})$ vanish. Numerically, this is equivalent to diagonalizing a 2$\times$2 matrix: $\xi_{i}(T_{0},n_{1}\mathbf{k}_{1})\sigma_{z}+\frac{\xi_{i}(T_{+},n_{1}\mathbf{k}_{1})+\xi_{i}(T_{-},n_{1}\mathbf{k}_{1})}{2}\sigma_{x}+\frac{\xi_{i}(T_{+},n_{1}\mathbf{k}_{1})-\xi_{i}(T_{-},n_{1}\mathbf{k}_{1})}{2}\sigma_{y}$ into $\tilde{\xi}_{i}(T_{0},n_{1}\mathbf{k}_{1})\sigma_{z}$. Note that the $\mathcal{T}$ symmetry enforces $\xi_{i}(T_{+},n_{1}\mathbf{k}_{1})=\xi_{i}^{*}(T_{-},n_{1}\mathbf{k}_{1})$, so this matrix is Hermitian and $\tilde{\xi}_{i}(T_{0},n_{1}\mathbf{k}_{1})\sigma_{z}$ is real. The constraint to $|\tau_{i}\rangle$ is $\tilde{\xi}_{i}(n_{m}\mathbf{k}_{j})=-e^{i\Theta_{n_{m}\mathbf{k}_{j}}}\tilde{\xi}_{i}(n_{m},-\mathbf{k}_{j})$, in which $\Theta_{n_{m}\mathbf{k}_{j}}$ is the angle between $\langle n_{m}\mathbf{k}_{j}+|\vec{S}|n_{m}\mathbf{k}_{j}+\rangle$ and $\langle n_{m},-\mathbf{k}_{j},+|\vec{S}|n_{m},-\mathbf{k}_{j},+\rangle$ after the same rotations deciding $\tilde{\xi}_{i}(T_{0},n_{m}\mathbf{k}_{j})$ and $\tilde{\xi}_{i}(T_{0},n_{m},-\mathbf{k}_{j})$. The $\mathcal{T}$ symmetry enforces $\Theta_{n_{m}\mathbf{k}_{j}}=\pm\pi$.
@@
Materials Energy cutoff (Ha) QE $k$-mesh QE $q$-mesh EPW $k$-mesh EPW $q$-mesh
Bi2Se3 (slab) 60 30$\times$30$\times$1 4$\times$4$\times$1 180$\times$180$\times$1 180$\times$180$\times$1
Bi2Se3 40 12$\times$12$\times$12 3$\times$3$\times$3 20$\times$20$\times$20 20$\times$20$\times$20
SnTe 60 12$\times$12$\times$12 6$\times$6$\times$6 20$\times$20$\times$20 20$\times$20$\times$20
Bi2Pd 60 12$\times$12$\times$12 4$\times$4$\times$4 20$\times$20$\times$20 20$\times$20$\times$20
Al 80 10$\times$10$\times$10 10$\times$10$\times$10 30$\times$30$\times$30 30$\times$30$\times$30
Pd 60 12$\times$12$\times$12 6$\times$6$\times$6 20$\times$20$\times$20 20$\times$20$\times$20
Cd 60 12$\times$12$\times$12 6$\times$6$\times$6 20$\times$20$\times$20 20$\times$20$\times$20
Ta 60 10$\times$10$\times$10 10$\times$10$\times$10 20$\times$20$\times$20 20$\times$20$\times$20
Pb 60 12$\times$12$\times$12 6$\times$6$\times$6 20$\times$20$\times$20 20$\times$20$\times$20
Hg 70 18$\times$18$\times$18 3$\times$3$\times$3 18$\times$18$\times$18 18$\times$18$\times$18
@@ III Applications
Below we apply DDM to several representative materials. The input data to construct $\mathbb{V}$ are obtained from Quantum Espresso with GGA-PBE exchange correlation functional [16, 17, 18], optimized norm-conserving Vanderbilt pseudopotential [19, 20] and a 0.01 Ry Methfessel-Paxton smearing [21] for all materials. The convergence criteria are $10^{-8}(a,u)$ for lattice and atomic relaxations, and $10^{-12}(a,u)$ for electronic iterations. The EPCs from Quantum Espresso are further Wannier interpolated by using the EPW code [22]. Material-dependent setups are summarized in Tab. I. Our DDM code to analyze $\mathbb{V}$ is accessible at: [https://github.com/ZimengZeng-THU/Resolving-different-superconducting-pairing].
@@ III.1 Bi2Se3 slab
Figure 1: (a) Atomic structure of a two-quintuple layer Bi2Se3 slab employed for the DFT+DFPT calculation; (b) Electron band structure; the shadowed region indicates the range of the electron-doped Fermi level for the EPC and DDM calculations; (c) Leading pairing strength in the spin-singlet and spin-triplet sectors; the insets schematically decompose the Fermi surfaces into two circles localized on the two surfaces of the slabs. The symmetry of the corresponding DDM eigenvector [c.f. Eq. (16)] is reflected by plotting the signs of its components at a selected pair of ±$\pm\mathbf{k}$ points.
We first consider an intuitively transparent example - the well-known 3D toplogical insulator Bi2Se3 [23]. By cleaving a slab [Fig. 1(a)], helical bands form on both surfaces [Fig. 1(b)]. Within the bulk gap, the helical surface states lead to a doubly-degenerate FS. Since the other bulk bands are fully gaped, the band index $n$ can be omitted. According to our definitions [Eqs. (5) and (9)], $\hat{\Delta}(\mathbf{k}\pm\mp)$ refers to intra-surface pairing, and $\hat{\Delta}(\mathbf{k}\pm\pm)$ refers to inter-surface pairing.
When the slab is infinitely thick, the two surfaces are decoupled. Therefore, $\langle\hat{\Delta}(\mathbf{k}\pm\pm)\rangle=0$, and there is no constraint on the relative phase of $\langle\hat{\Delta}(\mathbf{k}+-)\rangle$ and $\langle\hat{\Delta}(\mathbf{k}-+)\rangle$, indicating degeneracy between the leading $\langle\hat{\Delta}^{S}\rangle$ and $\langle\hat{\Delta}^{T_{0}}\rangle$. For a finite thick slab, Cooper pairs related by spin-conserving inter-surface scatterings, e.g. $\tilde{c}_{\mathbf{-k},+}c_{\mathbf{k},-}\rightarrow\tilde{c}_{\mathbf{k},+}c_{\mathbf{-k},-}$, tend to establish a unified phase to take advantage of the attraction, making $\Delta^{S}$ the dominating pairing channel. This argument suggests that the spatial separation of the surface states makes the the leading pairing channels from the odd-parity and even-parity sectors close in strength.
@@ III.2 Doped Bi2Se3 and SnTe
Figure 3: DDM analysis results of electron-doped SeTe. The eigenvalues ($\lambda_{i}$) and eigenvectors are classified by the irreps. of the Oh point group. The color contour shows the sign and magnitude of the eigenvector components on the Fermi surface.
@@
For doped Bi2Se3, the leading even-parity pairing $\lambda_{A_{1g}}$ is calculated to be 0.315, in good agreement with the isotropic $\lambda$ reported in [28]. The leading odd-parity pairing $\lambda_{Eu}$ is one order of magnitude smaller than $\lambda_{A_{1g}}$, in agreement with Namoto et al.’s qualitative conclusion based on the weak-SOC approximation [12]. We should mention that Wan and Savarasov [6] in contrast predicted large odd-parity pairings $\lambda_{A_{2u}}$ and $\lambda_{E_{u}}$ based on the FS harmonics projection. This prediction was rationalized by a singular EPC associated with the acoustic phonon at $q$=(0,0,0.04) (in terms of the reciprocal vectors). However, this singular EPC was not reproduced in either [28] or [12]. We note that our phonon linewidth calculation gives very similar results to those reported in [28] and [12]. In particular, no singularity is observed at $q$=(0,0,0.04).
With respect to doped SnTe, DDM predicts that there is no odd-parity pairing channel of the same order of magnitude in strength as the leading even-parity pairing channel. The ratio between $\lambda_{A_{1u}/E_{u}}$ and $\lambda_{A_{1g}}$ is roughly around 0.1, in agreement with [12]. The leading odd-parity gap function shown in [12] can also be nicely compared with our Eu(1) eigenstate as plotted in Fig. 3. The other odd-pairing eigenstates of comparative strength have not been shown before.
@@ III.3 Elemental metals
Figure 4: leading even-parity(black),leading odd-parity(blue),sub-leading even-parity(red) superconducting pairing channels are calculated by our method
@@
Figure 4 summarizes $\lambda^{S(T)}_{max}$ of superconducting elemental metals, including Pb,Ta,Al,Pd,Hg,Cd. For all these materials, $\lambda^{S}_{max}$ obtained from the DDM method can be well compared with the isotropic dimensionless $\lambda$ reported in literature [29, 30, 31]. The ratio $\lambda_{max}^{T}/\lambda_{max}^{S}$ varies between 0.1 and 0.5. It is interesting to note that the highest ratio occurs in Pb and Hg, in which EPC is known to be quite large. In experiment, Pb is a frequently used component to search for topological superconductivity [32, 33, 34, 35]. In Fig. 5, we plot the leading pairing channels of Pb resolved by DDM. It is worth further investigations to identify the key factors for a high $\lambda_{max}^{T}/\lambda_{max}^{S}$ ratio.

### v3

@@ II.3 DDM solution
$|\sigma_{i}\rangle$ $=$ $[\xi_{i}(n_{1},\mathbf{k}_{1}),\xi_{i}(n_{1},\mathbf{k}_{2}),...]^{T},$ (16)
$|\tau_{i}\rangle$ $=$ $[\xi_{i}(T_{0},n_{1},\mathbf{k}_{1}),\xi_{i}(T_{+},n_{1},\mathbf{k}_{1}),\xi_{i}(T_{-},n_{1},\mathbf{k}_{1}),...]^{T}$
reflect the $\mathbf{k}$-space symmetry of the corresponding pairing channel, which can be classified by the irreducible representations (irreps) of the symmetry group.
There is still one remaining issue. So far, $\hat{\Delta}(n,\mathbf{k},s_{1},s_{2})=\tilde{c}_{n,\mathbf{-k},s_{2}}c_{n,\mathbf{k},s_{1}}$ and $\hat{\Delta}(n,-\mathbf{k},s_{2},s_{1})=\tilde{c}_{n,\mathbf{k},s_{1}}c_{n,\mathbf{-k},s_{2}}$ have been treated as independent operators, which are actually related by an undecided gauge transformation and the fermion anti-commutation relation. (Mathematically, the basis of $\mathbb{V}$ is overcomplete.) Therefore, additional constraints should be applied to the eigenvectors.
Specifically, the symmetrized even-parity operator $\hat{\Delta}^{S}(n,\mathbf{k})$ is U(1) and SU(2) invariant. The constraint to $|\sigma_{i}\rangle$ is simply $\xi_{i}(n_{m},\mathbf{k}_{j})=\xi_{i}(n_{m},-\mathbf{k}_{j})$. In practice, it means that eigenvectors inconsistent with it should be discarded.
For the odd-parity sector, the eigenvector coefficients $\xi_{i}(T_{0},n_{1},\mathbf{k}_{1})$, $\xi_{i}(T_{+},n_{1},\mathbf{k}_{1})$ and $\xi_{i}(T_{-},n_{1},\mathbf{k}_{1})$ mutually transform under the SU(2) rotation like a spin-1 object. Our recipe is to find a rotation such that $\xi_{i}(T_{+},n_{1},\mathbf{k}_{1})$ and $\xi_{i}(T_{-},n_{1},\mathbf{k}_{1})$ vanish. Denote the eigenvector after the rotation as $|\tau_{i}\rangle=[\tilde{\xi}_{i}(T_{0},n_{1},k_{1}),0,0,\tilde{\xi}_{i}(T_{0},n_{2},k_{2}),0,0,...]^{T}$. The constraint to $|\tau_{i}\rangle$ is $\tilde{\xi}_{i}(n_{m},\mathbf{k}_{j})=-e^{i\Theta_{n_{m},\mathbf{k}_{j}}}\tilde{\xi}_{i}(n_{m},-\mathbf{k}_{j})$, in which $\Theta_{n_{m},\mathbf{k}_{j}}$ is the angle between $\langle n_{m},\mathbf{k}_{j},+|\vec{S}|n_{m},\mathbf{k}_{j},+\rangle$ and $\langle n_{m},-\mathbf{k}_{j},+|\vec{S}|n_{m},-\mathbf{k}_{j},+\rangle$ after the same rotations deciding $\tilde{\xi}_{i}(T_{0},n_{m},\mathbf{k}_{j})$ and $\tilde{\xi}_{i}(T_{0},n_{m},-\mathbf{k}_{j})$. The $\mathcal{T}$ symmetry enforces $\Theta_{n_{m},\mathbf{k}_{j}}=0$ or $\pi$.
@@
Materials Energy cutoff (Ry) Smearing(Ry) QE $k$-mesh QE $q$-mesh EPW $k$-mesh EPW $q$-mesh
Bi2Se3 (slab) 60 0.02 30$\times$30$\times$1 4$\times$4$\times$1 180$\times$180$\times$1 180$\times$180$\times$1
Bi2Se3 40 0.01 12$\times$12$\times$12 3$\times$3$\times$3 20$\times$20$\times$20 20$\times$20$\times$20
SnTe 60 0.02 12$\times$12$\times$12 6$\times$6$\times$6 20$\times$20$\times$20 20$\times$20$\times$20
Al 80 0.02 10$\times$10$\times$10 10$\times$10$\times$10 30$\times$30$\times$30 30$\times$30$\times$30
Pd 60 0.02 12$\times$12$\times$12 6$\times$6$\times$6 20$\times$20$\times$20 20$\times$20$\times$20
Cd 60 0.02 12$\times$12$\times$12 6$\times$6$\times$6 20$\times$20$\times$20 20$\times$20$\times$20
Ta 60 0.02 10$\times$10$\times$10 10$\times$10$\times$10 20$\times$20$\times$20 20$\times$20$\times$20
Pb 90 0.01 12$\times$12$\times$12 6$\times$6$\times$6 20$\times$20$\times$20 20$\times$20$\times$20
Hg 70 0.02 18$\times$18$\times$18 3$\times$3$\times$3 18$\times$18$\times$18 18$\times$18$\times$18
@@ III Applications
Below we apply DDM to several representative materials. The input data to construct $\mathbb{V}$ are obtained from QUANTUM ESPRESSO with generalized gradient approximation Perdew-Burke-Ernzerhof exchange correlation functional [16, 17, 18], optimized norm-conserving Vanderbilt pseudopotential [19] and Methfessel-Paxton smearing [20] for all materials. The convergence criteria are $10^{-6}(a.u)$ for lattice and atomic relaxations, and $10^{-12}(a.u)$ for electronic iterations. The EPCs from QUANTUM ESPRESSO are further Wannier interpolated by using the EPW code [21]. Material-dependent setups are summarized in Table LABEL:m. Our DDM code to analyze $\mathbb{V}$ is accessible online [22].
@@ III.1 Bi2Se3 slab
We first consider an intuitively transparent example - the well-known 3D topological insulator Bi2Se3 [23]. By cleaving a slab [Fig. 1(a)], helical bands form on both surfaces [Fig. 1(b)]. Within the bulk gap, the helical surface states lead to a doubly-degenerate FS. Since the other bulk bands are fully gapped, the band index $n$ can be omitted. According to our definitions [Eqs. (5) and (9)], $\hat{\Delta}(\mathbf{k},\pm,\mp)$ refers to intra-surface pairing, and $\hat{\Delta}(\mathbf{k},\pm,\pm)$ refers to inter-surface pairing.
When the slab is infinitely thick, the two surfaces are decoupled. Therefore, $\langle\hat{\Delta}(\mathbf{k},\pm,\pm)\rangle=0$, and there is no constraint on the relative phase of $\langle\hat{\Delta}(\mathbf{k},+,-)\rangle$ and $\langle\hat{\Delta}(\mathbf{k},-,+)\rangle$, indicating degeneracy between the leading $\langle\hat{\Delta}^{S}\rangle$ and $\langle\hat{\Delta}^{T_{0}}\rangle$. For a finite thick slab, Cooper pairs related by spin-conserving inter-surface scatterings, e.g., $\tilde{c}_{\mathbf{-k},+}c_{\mathbf{k},-}\rightarrow\tilde{c}_{\mathbf{k},+}c_{\mathbf{-k},-}$, tend to establish a unified phase to take advantage of the attraction, making $\Delta^{S}$ the dominating pairing channel. This argument suggests that the spatial separation of the surface states makes the leading pairing channels from the odd-parity and even-parity sectors close in strength.
@@
Figure 1: (a) Atomic structure of a two-quintuple layer Bi2Se3 slab employed for the DFT+DFPT calculation; (b) Electron band structure; the shadowed region indicates the range of the electron-doped Fermi level(the bule dash line) for the EPC and DDM calculations. The red dash line means undoped fermi level; (c) Leading pairing strength in the spin-singlet and spin-triplet sectors; the insets schematically decompose the Fermi surfaces into two circles localized on the two surfaces of the slabs. The symmetry of the corresponding DDM eigenvector [cf. Eq. (16)] is reflected by plotting the signs of its components at a selected pair of ±$\pm\mathbf{k}$ points.
@@ III.2 Doped Bi2Se3 and SnTe
For doped Bi2Se3, the leading even-parity pairing $\lambda_{A_{1g}}$ is calculated to be 0.315, in good agreement with the isotropic $\lambda$ reported in Ref. [28]. The leading odd-parity pairing $\lambda_{Eu}$ is one order of magnitude smaller than $\lambda_{A_{1g}}$, in agreement with Nomoto $et$ $al.$’s qualitative conclusion based on the weak-SOC approximation [12]. We should mention that Wan and Savrasov [6] in contrast predicted large odd-parity pairings $\lambda_{A_{2u}}$ and $\lambda_{E_{u}}$ based on the FS harmonics projection. This prediction was rationalized by a singular EPC associated with the acoustic phonon at $q$=(0,0,0.04) (in terms of the reciprocal vectors). However, this singular EPC was not reproduced in either Ref. [28] or [12]. We note that our phonon linewidth calculation gives very similar results to those reported in [28] and [12]. In particular, no singularity is observed at $q$=(0,0,0.04).
With respect to doped SnTe, DDM predicts that there is no odd-parity pairing channel of the same order of magnitude in strength as the leading even-parity pairing channel. The ratio between $\lambda_{A_{1u}/E_{u}}$ and $\lambda_{A_{1g}}$ is roughly around 0.1, in agreement with Ref. [12]. The leading odd-parity gap function shown in Ref. [12] can also be nicely compared with our Eu(1) eigenstate as plotted in Fig. 3. The other odd-pairing eigenstates of comparative strength have not been shown before.
Figure 3: DDM analysis results of electron-doped SnTe. The eigenvalues ($\lambda_{i}$) and eigenvectors are classified by the irreps. of the Oh point group. The color contour shows the sign and magnitude of the eigenvector components on the Fermi surface.
@@ III.3 Elemental metals
Figure 4: $\lambda_{max}^{S}$ (grey bar), $\lambda_{max}^{T}$ (blue bar), and their ratios (red circles) of elemental metals.
@@
Figure 6: $A_{1u}$ eigenvector of Pb in the d-vector representation: (a) d-vectors on the FS are plotted as arrows; (b)-(d) plot the three components of d-vector, respectively.
Figure 4 summarizes $\lambda^{S(T)}_{max}$ of superconducting elemental metals, including Pb,Ta,Al,Pd,Hg,Cd. For all these materials, $\lambda^{S}_{max}$ obtained from the DDM can be well compared with the isotropic dimensionless $\lambda$ reported in literature [29, 30, 31]. The ratio $\lambda_{max}^{T}/\lambda_{max}^{S}$ varies between 0.1 and 0.5. It is interesting to note that the highest ratio occurs in Pb and Hg, in which EPC is known to be quite large. In experiment, Pb is a frequently used component to synthesize potential topological superconductor candidates [32, 33, 34, 35].
Introducing the effect of Coulomb repulsion in general suppresses spin-singlet pairing, making spin-triplet pairing more favorable. We have tested subtracting a constant $\mu^{*}$(renormalized by a density of states factor) from Eq. (8). We approximate the Coulomb repulsion as momentum independent but pseudospin conserving, i.e., modifying pairing matrix elements with $s_{1}=s_{3}$ and $s_{2}=s_{4}$ only. Such a real-space delta repulsion effectively suppresses $\lambda_{max}^{S}$ while $\lambda_{max}^{T}$ is nearly unperturbed. Within this treatment, the critical $\mu^{*}$ to make $\lambda_{max}^{T}$ dominate is 0.7 for Pb, and 0.82 for Hg, very close to $\lambda_{max}^{S}$-$\lambda_{max}^{T}$.
The realistic critical $\mu^{*}$ could be much lower, because we know that in McMillan’s empirical formula for the conventional singlet pairing, $T_{c}^{S}\propto\exp\left[-\frac{1.04(1+\lambda)}{\lambda-\mu^{*}(1+0.62\lambda)}\right]$ [3]. The effect of $\mu^{*}$ is enhanced by a factor of $(1+0.62\lambda)$. We do not have an empirical formula for the spin-triplet $T_{c}$ ($T_{c}^{T}$). If we naively substitute $\lambda_{max}^{S}$ and $\lambda_{max}^{T}$ into the same formula and consider $\mu^{S*}=\mu^{*}$ (effective in leading singlet channel), $\mu^{T*}=0$[36, 6] (ineffectice in triplet channel). The critical $\mu^{*}$ making $T_{c}^{T}>T_{c}^{S}$ reduces to 0.27 for Pb, and 0.22 for Hg. These values appear more accessible, but we should warn that they are still very crude estimations.
The DDM eigenvectors provide significant physical insights into the associating superconducting states, particularly for understanding the response to an external magnetic field. In Fig. 5, we plot the leading pairing channel of Pb. It is possible to transform these eigenvectors into the d-vector representation commonly used in theoretical literatures [37]. To do so, we apply an additional SU(2) transformation to rotate the spin expectation value of the Kohn-Sham wavefunctions to align a fixed axis. The same SU(2) transformation rotates the phase-fixed $\xi(T_{0},n,\mathbf{k})\sigma_{z}$ into a non-diagonal 2x2 matrix:
$\hat{\xi}\textquoteright_{i}(n,\mathbf{k})=\begin{pmatrix}\xi\textquoteright_{i,\uparrow^{\prime},\downarrow^{\prime}}(n,\mathbf{k})&\xi^{\prime}_{i,\uparrow^{\prime},\uparrow^{\prime}}(n,\mathbf{k})\\ \xi^{\prime}_{i,\downarrow^{\prime},\downarrow^{\prime}}(n,\mathbf{k})&\xi^{\prime}_{i,\downarrow^{\prime},\uparrow^{\prime}}(n,\mathbf{k})\end{pmatrix},$ (17)
in which the pseudo spin labels $\uparrow^{\prime},\downarrow^{\prime}$, have been globally aligned. The d-vector is then defined as
$\mathbf{d}(n,\mathbf{k})=\begin{pmatrix}\frac{\xi^{\prime}_{i,\uparrow^{\prime},\uparrow^{\prime}}(n,\mathbf{k})+\xi^{\prime}_{i,\downarrow^{\prime},\downarrow^{\prime}}(n,\mathbf{k})}{\sqrt{2}}\\ \frac{i\xi^{\prime}_{i,\uparrow^{\prime},\uparrow^{\prime}}(n,\mathbf{k})-i\xi^{\prime}_{i,\downarrow^{\prime},\downarrow^{\prime}}(n,\mathbf{k})}{\sqrt{2}}\\ \frac{\xi^{\prime}_{i,\uparrow^{\prime},\downarrow^{\prime}}(n,\mathbf{k})-\xi^{\prime}_{i,\downarrow^{\prime},\uparrow^{\prime}}(n,\mathbf{k})}{\sqrt{2}}\end{pmatrix}.$ (18)
Using the leading spin-triplet channel ($A_{1u}$) of Pb as an example, we make the vector plot of $\mathbf{d}(n,\mathbf{k})$, as well as the color plots of each component in Fig. 6.